RELATIONSHIPS FOR MOMENTS OF ORDER STATISTICS FROM THE RIGHT-TRUNCATED GENERALIZED HALF

LOGISTIC DISTRIBUTION

N. BALAKRISHNAN AND RITA AGGARWALA

Department of Mathematics and Statistics, McMaster University, 1280 Main Street West, Hamilton, Ontario, Canada L8S 4K1

(Received February 13, 1995; revised June 2, 1995)

Abstract. In this paper, we establish several recurrence relations satisfied by the single and the product moments for order statistics from the righttruncated generalized half logistic distribution. These relationships may be used in a simple recursive manner in order to compute the single and the product moments of all order statistics for all sample sizes and for any choice of the truncation parameter P. These generalize the corresponding results for the generalized half logistic distribution derived recently by Balakrishnan and Sandhu (1995, J. Statist. Comput. Simulation, 52, 385–398).

Key words and phrases: Order statistics, recurrence relations, single moments, product moments, generalized half logistic distribution, right-truncated distribution, IFR family.